# Bilevel Reinforcement Learning via the Development of Hyper-gradient without Lower-Level Convexity

### v1

@@ 3.2 Bilevel reinforcement learning formulation
where the upper-level function $f$ is defined on $\mathbb{R}^{n}\times\mathbb{R}^{\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|}$.
3.3 Applications of bilevel reinforcement learning
We introduce two examples unified in the formulation (4). To make a distinction between environments at two levels, we denote the upper-level MDP by ¯$\bar{\mathcal{M}}_{\bar{\tau}}=\{\mathcal{S},\mathcal{A},\bar{P},\bar{r},\bar{\gamma},\bar{\tau}\}$ with an initial state distribution $\bar{\rho}$, and the lower-level MDP by $\mathcal{M}_{\tau}(x)$ with $\rho$.
Reward shaping. In reinforcement learning, the reward function acts as the guiding signal to motivate agents to achieve specified goals. However, in many cases, the rewards are sparse, which impedes the policy learning, or are partially incorrect, which leads to inaccurate policies. To this end, from the perspective of bilevel RL, it is advisable to shape an auxiliary reward function $r(x)$ at the lower level for efficient agent training, while maintaining the original environment at the upper level to align with the initial task evaluation [34], which is established as
@@ 3.3 Applications of bilevel reinforcement learning
Reinforcement learning from human feedback (RLHF). The target of RLHF is to learn the intrinsic reward function that incorporates expert knowledge, from simple labels only containing human preferences. Drawing from the original framework [14], Chakraborty et al. [10] and Shen et al. [69] have formulated it in a bilevel form, which optimizes a policy under the parameterized $r(x)$ at the lower level, and adjusts $x$ to align the preference predicted by the reward model $r(x)$ with the true labels at the upper level.
$\min\limits_{x\in\mathbb{R}^{n}}\mathbb{E}_{y,d_{1},d_{2}\sim\rho\left(d;\pi^{*}\left(x\right)\right)}\left[l_{h}\left(d_{1},d_{2},y;x\right)\right],\quad\mathrm{s.\,t.}\ \pi^{*}\left(x\right)=\operatorname*{arg\,min}\limits_{\pi\in\Delta^{\left|{\mathcal{A}}\right|}}-V_{\mathcal{M}_{\tau}\left(x\right)}^{\pi}\left(\rho\right),$
@@
$P(d_{i})=\bar{\rho}\left({s_{0}^{i}}\right)\left[{\Pi_{h=0}^{H-2}\pi^{*}(x)\left({a_{h}^{i}|s_{h}^{i}}\right)\bar{P}\left({s_{h+1}^{i}|s_{h}^{i},a_{h}^{i}}\right)}\right]\pi^{*}(x)\left({a_{H-1}^{i}|s_{H-1}^{i}}\right),$
and the preference label $y\in\{0,1\}$, indicating preference for $d_{1}$ over $d_{2}$, obeys human feedback distribution $y\sim D_{human}(y|d_{1},d_{2})$. Moreover, $l_{h}$ is the binary cross-entropy loss,
$l_{h}\left(d_{1},d_{2},y;x\right)=-y\log P\left(d_{1}\succ d_{2};x\right)-\left(1-y\right)\log P\left(d_{2}\succ d_{1};x\right),$
with the preference probability $P\left({d_{1}\succ d_{2};x}\right)$ built by the Bradley–Terry model:
$P\left(d_{1}\succ d_{2};x\right)=\frac{\exp\left(\sum_{h=0}^{H-1}{r_{s_{h}^{1},a_{h}^{1}}\left(x\right)}\right)}{\exp\left(\sum_{h=0}^{H-1}{r_{s_{h}^{1},a_{h}^{1}}\left(x\right)}\right)+\exp\left(\sum_{h=0}^{H-1}{r_{s_{h}^{2},a_{h}^{2}}\left(x\right)}\right)}.$
@@ 4 Model-Based Soft Bilevel Reinforcement Learning
Recall that the optimal soft quantities adhere to the following softmax temporal value consistency conditions [62]:
$\forall(s,a)\in\mathcal{S}\times\mathcal{A}:\,\,\,\,\,\ \pi_{sa}^{*}(x)$ $=\exp\left(\tau^{-1}\left({Q_{sa}^{*}\left(x\right)-V_{s}^{*}\left(x\right)}\right)\right)=\frac{\exp\left\{Q_{sa}^{*}\left(x\right)/\tau\right\}}{\sum_{a^{\prime}}{\exp\left(Q_{sa^{\prime}}^{*}\left(x\right)/\tau\right)}},$ (5)
$\forall s\in\mathcal{S}:\quad V_{s}^{*}\left(x\right)$ $=\tau\log\left(\sum\nolimits_{a}{\exp\left(\frac{r_{sa}\left(x\right)+\gamma\sum_{s^{\prime}}{P_{sas^{\prime}}V_{s^{\prime}}^{*}\left(x\right)}}{\tau}\right)}\right)$ (6)
@@
$|\displaystyle\varphi(x,v):\ \ \mathbb{R}^{n}\times\mathbb{R}^{\left|{\mathcal{S}}\right|}$$\longrightarrow\mathbb{R}^{\left|{\mathcal{S}}\right|}$
@@
$\nabla_{v}\varphi\left({x,V^{*}(x)}\right)=\gamma P^{\pi^{*}(x)}.$
Combining the above proposition with (8), we identify the hyper-gradient $\nabla\phi(x)$ related to the problem (4),
$-\tau^{-1}\nabla_{x}\varphi\left({x,V^{*}(x)}\right)^{\top}\left({I-\gamma P^{\pi^{*}(x)}}\right)^{-\top}U^{\top}\operatorname{diag}\left({\pi^{*}(x)}\right)\nabla_{\pi}f(x,\pi^{*}(x)).$ (9)
In this manner, we unveil the hyper-gradient in the context of bilevel RL, by harnessing the softmax temporal value consistency and derivatives of a fixed-point equation.
0:  outer and inner iteration numbers $K,N$, step sizes $\beta,\xi$, initialization $x_{0},\pi_{0},Q_{0},w_{0}$
1:  for $k=1,\dots,K$ do
2:     $A_{k}=\left({I-\gamma P^{\pi_{k}}}\right)^{\top},\quad b_{k}=U^{\top}\operatorname{diag}\left({\pi_{k}}\right)\nabla_{\pi}f(x_{k},\pi_{k})$
3:     Compute $V_{k}$ by (10)
4:     $w_{k}=w_{k-1}-\xi\left({\left({A_{k}^{\top}A_{k}}\right)w_{k-1}-A_{k}^{\top}b_{k}}\right)$
6:     $Q_{k,0}=Q_{k}$
7:     for $n=1,\dots,N$ do
8:        $Q_{k,n}=\mathcal{T}_{\mathcal{M}_{\tau}(x_{k+1})}(Q_{k,n-1})$
9:     end for
11:  end for
11:  $(x_{K+1},\pi_{K})$
In light of the characterization (9), a model-based soft bilevel reinforcement learning algorithm, called M-SoBiRL, is proposed in Algorithm 1. In summary, the $k$-th outer iteration performs an inexact hyper-gradient descent step on $x_{k}$, with the aid of auxiliary iterates $\left({\pi_{k},Q_{k},V_{k},w_{k}}\right)$. Specifically, given $x_{k+1}$, the inner iterations aim to (approximately) solve the lower-level entropy-regularized MDP problem. To this end, the soft policy iteration studied by [29, 9] inspires us to define the following soft Bellman optimality operator $\mathcal{T}_{\mathcal{M}_{\tau}\left(x\right)}:\mathbb{R}^{\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|}\to\mathbb{R}^{\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|}$ associated with $x\in\mathbb{R}^{n}$,
$\forall\ (s,a)\in\mathcal{S}\times\mathcal{A}:\quad\mathcal{T}_{\mathcal{M}_{\tau}\left(x\right)}(Q)(s,a)=r_{sa}(x)+\gamma\underset{s^{\prime}\sim P(\cdot\mid s,a)}{\mathbb{E}}\left[\tau\log\left(\left\|\exp\left(Q\left(s^{\prime},\cdot\right)/\tau\right)\right\|_{1}\right)\right].$
Applying this operator iteratively leads to the optimal soft Q-value function, with a linear convergence rate in theory (see Section G.4). Therefore, $N$ inner iterations are invoked in line 7-9 of Algorithm 1 to estimate $Q^{*}\left({x_{k+1}}\right)$, and the warm-start strategy is adopted in line 10 of Algorithm 1 to initialize the next outer iteration with historical information, where it amortizes the computation of $Q^{*}(x)$ through the outer iterations. Additionally, in order to recover $\pi_{k+1}$ from $Q_{k+1}$, we take into account the consistency condition (5) which involves the softmax mapping as follows,
@@
$\forall s\in\mathcal{S}:\quad V_{k}\left(s\right)=\tau\log\left(\sum\nolimits_{a}{\exp\left(Q_{k}\left(s,a\right)/\tau\right)}\right),$ (10)
which complies with the form of (7). Finally, collecting the iterates $\left(x_{k},\pi_{k},V_{k},w_{k}\right)$ yields the model-based hyper-gradient estimator ^$\widehat{\nabla}\phi$ in line 5 of Algorithm 1, i.e.,
^$\widehat{\nabla}\phi\left(x_{k},\pi_{k},V_{k},w_{k}\right):=\nabla_{x}f\left(x_{k},\pi_{k}\right)+\frac{1}{\tau}{\nabla r(x_{k})}^{\top}\operatorname{diag}\left({\pi_{k}}\right)\nabla_{\pi}f(x_{k},\pi_{k})-\frac{1}{\tau}\nabla_{x}\varphi(x_{k},V_{k})^{\top}w_{k}.$
@@ 5 Model-Free Soft Bilevel Reinforcement Learning
In many real-world applications, agents lack access to the accurate model $P$ of the environment, underscoring the need for efficient model-free algorithms [65, 66, 46, 61, 67, 29]. When $P$ is unknown, two obstacles appear from the hyper-gradient computation (9). Specifically, it explicitly relies on the black-box transition matrix $P$ to compute $\nabla V^{*}(x)$, i.e.,
@@
and the computation (9) involves complicated large-scale matrix multiplications. To circumvent them, we demonstrate how to derive the hyper-gradient by an expectation, which allows us to estimate $\nabla\phi(x)$ via sampling fully first-order information. Following these developments, we offer a model-free soft bilevel reinforcement learning algorithm.
@@
$f\left(x,\pi\right)=\mathbb{E}_{d_{i}\sim\rho\left(d;\pi\right)}\left[l\left(d_{1},d_{2},...,d_{I};x\right)\right],$ (12)
where each trajectory $d_{i}=\{\left({s_{h}^{i},a_{h}^{i}}\right)\}_{h=0}^{H-1}$ ($i=1,2,\ldots,I$) is sampled from the trajectory distribution $\rho\left({d;\pi^{*}(x)}\right)$ generated by the policy $\pi^{*}(x)$ in the upper-level ¯$\bar{\mathcal{M}}_{\bar{\tau}}$, and $l(\cdot;x)$ associated with $x$ is a function of trajectories. Notice that it incorporates the applications in Section 3.3, with $I=1$, $H=1$, $l=Q^{\pi^{*}(x)}_{\bar{\mathcal{M}}_{\bar{\tau}}}\left({s_{0}^{1},a_{0}^{1}}\right)$ for reward shaping, and $I=2$, finite $H$, $l=\mathbb{E}_{y}\left[{l_{h}\left({d_{1},d_{2},y;x}\right)}\right]$ for RLHF.
@@
Nevertheless, it is still intractable to construct the ${\left|{\mathcal{S}}\right|\times n}$ matrix $\nabla V^{*}(x)$ based on the above element-wise calculation, not to mention the large-scale matrix multiplications in (9). To bypass these matrix computations, we resort to the “log probability trick” [75] and the consistency condition (5). The subsequent proposition confirms that we can evaluate the hyper-gradient by interacting with the environment and collecting fully first-order information.
@@
$\nabla\phi(x)=$ $\ \mathbb{E}_{d_{i}\sim\rho\left(d;\pi^{*}\left(x\right)\right)}\left[\nabla l\left(d_{1},d_{2},...,d_{I};x\right)\right]$ (13)
$+\tau^{-1}\mathbb{E}_{d_{i}\sim\rho\left(d;\pi^{*}\left(x\right)\right)}\left[l\left(d_{1},d_{2},...,d_{I};x\right)\left(\sum_{i}{\sum_{h}{\nabla\left(Q_{s_{h}^{i}a_{h}^{i}}^{*}\left(x\right)-V_{s_{h}^{i}}^{*}\left(x\right)\right)}}\right)\right].$ (14)
Essentially, the first term (13) in the hyper-gradient contributes to decreasing the upper-level function, and the second term (14) condenses the gradient information transmitted from the lower level. Aggregating these two directions constructs the hyper-gradient, and updating the upper-level variable $x$ along $-\nabla\phi(x)$ will lead to a first-order stationary point (see Theorem 6.6).
@@
~$\widetilde{\nabla}Q_{sa}(x_{k},\pi_{k})$ $=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}\nabla r_{s_{t},a_{t}}(x_{k})\ \Big{|}\ s_{0}=s,a_{0}=a,\pi_{k},\mathcal{M}_{\tau}\left({x_{k}}\right)\right];$
@@
~$\widetilde{\nabla}\phi(x_{k},\pi_{k})=$ $\ \mathbb{E}_{d_{i}\sim\rho\left(d;\pi_{k}\right)}\left[\nabla l\left(d_{1},d_{2},...,d_{I};x_{k}\right)\right]$ (15)
$+\tau^{-1}\mathbb{E}_{d_{i}\sim\rho\left(d;\pi_{k}\right)}\left[l\left(d_{1},d_{2},...,d_{I};x_{k}\right)\left(\sum_{i}{\sum_{h}{\widetilde{\nabla}\left(Q_{s_{h}^{i}a_{h}^{i}}-V_{s_{h}^{i}}\right)\left({x_{k},\pi_{k}}\right)}}\right)\right].$
Consequently, we propose a model-free soft bilevel reinforcement learning algorithm, called SoBiRL, which is outlined in Algorithm 2. In the $k$-th outer iteration, we search for an approximate optimal soft policy $\pi_{k}$ satisfying ${\left\|{{\pi}_{k}-\pi^{*}(x_{k})}\right\|^{2}_{2}\leq\epsilon}$, which can be achieved by executing $\mathcal{O}\left({\log\epsilon^{-1}}\right)$ iterations of the policy mirror descent algorithm [41, 84]. Then, we utilize ~$\widetilde{\nabla}\phi(x_{k},\pi_{k})$, whose estimation only involves first-order oracles, for an inexact hyper-gradient descent step.
@@
2:     Solve the lower-level problem approximately with the initial guess $\pi_{k-1}$ to get ${\pi}_{k}$ such that $\left\|{{\pi}_{k}-\pi^{*}(x_{k})}\right\|_{2}^{2}\leq\epsilon$
3:     Compute the approximate hyper-gradient ~$\widetilde{\nabla}\phi(x_{k},\pi_{k})$

### v2

@@ 3.2 Bilevel RL Formulation
where the upper-level $f$ is defined on $\mathbb{R}^{n}\times\mathbb{R}^{\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|}$.
3.3 Applications of Bilevel RL
We introduce two examples unified in the formulation (4). To make a distinction between environments at two levels, we denote the upper-level MDP by ¯$\bar{\mathcal{M}}_{\bar{\tau}}=\{\mathcal{S},\mathcal{A},\bar{P},\bar{r},\bar{\gamma},\bar{\tau}\}$ with an initial state distribution $\bar{\rho}$, and the lower-level MDP by $\mathcal{M}_{\tau}(x)$ with $\rho$. More applications can be found in Appendix K.
Reward Shaping In reinforcement learning, the reward function acts as the guiding signal to motivate agents to achieve specified goals. However, in many cases, the rewards are sparse, which impedes the policy learning, or are partially incorrect, which leads to inaccurate policies. To this end, from the perspective of bilevel RL, it is advisable to shape an auxiliary reward function $r(x)$ at the lower level for efficient agent training, while maintaining the original environment at the upper level to align with the initial task evaluation (Hu et al.,, 2020), which is established as
@@ 3.3 Applications of Bilevel RL
Reinforcement Learning from Human Feedback (RLHF) The target of RLHF is to learn the intrinsic reward function that incorporates expert knowledge, from simple labels only containing human preferences. Drawing from the original framework (Christiano et al.,, 2017), Chakraborty et al., (2024) and Shen et al., (2024) have formulated it in a bilevel form, which optimizes a policy under $r(x)$ at the lower level, and adjusts $x$ to align the preference predicted by the reward model $r(x)$ with the true labels at the upper level.
$\begin{array}[]{cl}\min\limits_{x\in\mathbb{R}^{n}}&\mathbb{E}_{y,d_{1},d_{2}\sim\rho\left(d;\pi^{*}\left(x\right)\right)}\left[l_{h}\left(d_{1},d_{2},y;x\right)\right]\\ \mathrm{s.\,t.}&\pi^{*}\left(x\right)=\operatorname*{arg\,min}\limits_{\pi\in\Delta^{\left|{\mathcal{A}}\right|}}-V_{\mathcal{M}_{\tau}\left(x\right)}^{\pi}\left(\rho\right),\end{array}$
@@
$P(d_{i})=$ $\bar{\rho}\left({s_{0}^{i}}\right)\left[{\Pi_{h=0}^{H-2}\pi^{*}(x)\left({a_{h}^{i}|s_{h}^{i}}\right)\bar{P}\left({s_{h+1}^{i}|s_{h}^{i},a_{h}^{i}}\right)}\right]$
×$\times\pi^{*}(x)\left({a_{H-1}^{i}|s_{H-1}^{i}}\right),$
and the preference label $y\in\{0,1\}$, indicating preference for $d_{1}$ over $d_{2}$, obeys human feedback distribution $y\sim D_{human}(y|d_{1},d_{2})$. Moreover, $l_{h}$ is the binary cross-entropy loss, $l_{h}\left(d_{1},d_{2},y;x\right)=-y\log P\left(d_{1}\succ d_{2};x\right)-\left(1-y\right)\log P\left(d_{2}\succ d_{1};x\right)$, with the preference probability $P\left({d_{1}\succ d_{2};x}\right)$ built by the Bradley–Terry model; see (101) in the appendix.
@@ 4 MODEL-BASED SOFT BILEVEL REINFORCEMENT LEARNING
Recall that the optimal soft quantities adhere to the following softmax temporal value consistency conditions (Nachum et al.,, 2017): $\forall(s,a)\in\mathcal{S}\times\mathcal{A}$,
$\!\!\pi_{sa}^{*}(x)\!$ $=\!{\exp\left\{Q_{sa}^{*}\left(x\right)/\tau\right\}}\big{/}\big{(}{\sum_{a^{\prime}}{\exp\left(Q_{sa^{\prime}}^{*}\left(x\right)/\tau\right)}}\big{)},$ (5)
$\!\!V_{s}^{*}\left(x\right)\!$ $=\!\tau\log\!\left(\!\sum\limits_{a}{\exp\left(\frac{r_{sa}\left(x\right)+\gamma\sum_{s^{\prime}}{P_{sas^{\prime}}V_{s^{\prime}}^{*}\left(x\right)}}{\tau}\right)}\!\!\right)\!$ (6)
@@
$\varphi(x,v):\ \mathbb{R}^{n}\times\mathbb{R}^{\left|{\mathcal{S}}\right|}$ $\longrightarrow\mathbb{R}^{\left|{\mathcal{S}}\right|}$
@@
Additionally, $\nabla_{v}\varphi\left({x,V^{*}(x)}\right)$ coincides with the $\gamma$-scaled transition matrix induced by the optimal soft policy $\pi^{*}(x)$, i.e., $\nabla_{v}\varphi\left({x,V^{*}(x)}\right)=\gamma P^{\pi^{*}(x)}$.
Combining the above proposition with (8), we identify the hyper-gradient $\nabla\phi(x)$ to the problem (4); its expression is delayed in (24) of Appendix G. In this manner, we unveil the hyper-gradient in the context of bilevel RL, by harnessing the softmax temporal value consistency and derivatives of a fixed-point equation.
In light of the characterization of $\nabla\phi$, a model-based soft bilevel reinforcement learning algorithm, called M-SoBiRL, is proposed in Algorithm 2. In summary, the $k$-th outer iteration performs an inexact hyper-gradient descent step on $x_{k}$, with the aid of auxiliary iterates $\left({\pi_{k},Q_{k},V_{k},w_{k}}\right)$. Specifically, given $x_{k+1}$, the inner iterations aim to (approximately) solve the lower-level entropy-regularized MDP problem. To this end, the soft policy iteration studied by Haarnoja et al., 2018a ; Cen et al., (2022) inspires us to define the following soft Bellman optimality operator $\mathcal{T}_{\mathcal{M}_{\tau}\left(x\right)}:\mathbb{R}^{\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|}\to\mathbb{R}^{\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|}$ associated with $x\in\mathbb{R}^{n}$,
$\mathcal{T}_{\mathcal{M}_{\tau}\left(x\right)}(Q)(s,a):=r_{sa}(x)+\gamma{\mathbb{E}}\left[\tau\log\left(\left\|\exp\left(Q_{s^{\prime},\cdot}/\tau\right)\right\|_{1}\right)\right],$
where the expectation is taken with respect to $s^{\prime}\sim{P(\cdot\mid s,a)}$. Applying this operator iteratively leads to the optimal soft Q-value function, with a linear convergence rate in theory (see Section G.4). Therefore, $N$ inner iterations are invoked in line 7-9 of Algorithm 2 to estimate $Q^{*}\left({x_{k+1}}\right)$, and the warm-start strategy is adopted in line 10 of Algorithm 2 to initialize the next outer iteration with historical information, where it amortizes the computation of $Q^{*}(x)$ through the outer iterations. Additionally, in order to recover $\pi_{k+1}$ from $Q_{k+1}$, we take into account the consistency condition (5) which involves the softmax mapping as follows,
@@
$\!\forall s\in\mathcal{S}:\ V_{k}\left(s\right)=\tau\log\left(\sum\nolimits_{a}{\exp\left(Q_{k}\left(s,a\right)/\tau\right)}\right),\!$ (9)
which complies with the form of (7). Finally, collecting the iterates $\left(x_{k},\pi_{k},V_{k},w_{k}\right)$ yields the model-based hyper-gradient estimator ^$\widehat{\nabla}\phi$, i.e.,
^$\!\widehat{\nabla}\phi\left(x_{k},\pi_{k},V_{k},w_{k}\right):=$ $\,\nabla_{x}f\left(x_{k},\pi_{k}\right)\!+\!\frac{1}{\tau}{\nabla r(x_{k})}^{\top}\!\operatorname{diag}(\pi_{k})$
×$\times\!\nabla_{\pi}f(x_{k},\pi_{k})\!-\!\frac{1}{\tau}\nabla_{x}\varphi(x_{k},V_{k})^{\top}\!w_{k}.$
@@ 5 MODEL-FREE SOFT BILEVEL REINFORCEMENT LEARNING
In many real-world applications, agents lack access to the accurate model $P$ of the environment, underscoring the need for efficient model-free algorithms (Schulman et al.,, 2015, 2016; Lillicrap et al.,, 2016; Mnih et al.,, 2016; Schulman et al.,, 2017; Haarnoja et al., 2018a, ). When $P$ is unknown, two obstacles appear from the hyper-gradient computation (24). Specifically, it explicitly relies on the black-box transition matrix $P$ to compute $\nabla V^{*}(x)$, i.e.,
@@
and the computation involves complicated large-scale matrix multiplications. To circumvent them, we demonstrate how to derive the hyper-gradient in an expectation, which allows us to estimate $\nabla\phi(x)$ via sampling fully first-order information. Following these developments, we offer a model-free soft bilevel reinforcement learning algorithm.
@@
$f\left(x,\pi\right)=\mathbb{E}_{d_{i}\sim\rho\left(d;\pi\right)}\left[l\left(d_{1},d_{2},\ldots,d_{I};x\right)\right],$ (11)
where each trajectory $d_{i}=\{\left({s_{h}^{i},a_{h}^{i}}\right)\}_{h=0}^{H-1}$ is independently sampled from the trajectory distribution $\rho\left({d;\pi}\right)$ generated by the policy $\pi$ in the upper-level ¯$\bar{\mathcal{M}}_{\bar{\tau}}$, and $l(\cdot;x)$ associated with $x$ is a function of trajectories. Notice that it incorporates the applications in Section 3.3, with $I=1$, $H=1$, $l=-Q^{\pi^{*}(x)}_{\bar{\mathcal{M}}_{\bar{\tau}}}\left({s_{0}^{1},a_{0}^{1}}\right)$ for reward shaping, and $I=2$, finite $H$, $l=\mathbb{E}_{y}\left[{l_{h}\left({d_{1},d_{2},y;x}\right)}\right]$ for RLHF.
@@
$\nabla V_{s}^{*}(x)$ $=\mathbb{E}\big{[}\sum_{t=0}^{\infty}\gamma^{t}\nabla r_{s_{t},a_{t}}(x)\ \big{|}\ s_{0}=s,\pi^{*}(x)\big{]},$
$\nabla Q_{sa}^{*}(x)$ $=\mathbb{E}\big{[}\sum_{t=0}^{\infty}\gamma^{t}\nabla r_{s_{t},a_{t}}(x)\ \big{|}\ s_{0}\!=\!s,a_{0}\!=\!a,\pi^{*}(x)\big{]}.$
Nevertheless, it is still intractable to construct the ${\left|{\mathcal{S}}\right|\times n}$ matrix $\nabla V^{*}(x)$ based on the above element-wise calculation, not to mention the large-scale matrix multiplications in (24). To bypass these matrix computations, we resort to the “log probability trick” (Sutton and Barto,, 2018) and the consistency condition (5). The subsequent proposition confirms that we can evaluate the hyper-gradient by interacting with the environment and collecting fully first-order information.
@@
$\nabla\phi(x)=$ $\ \mathbb{E}_{d_{i}\sim\rho\left(d;\pi^{*}\left(x\right)\right)}\left[\nabla l\left(d_{1},d_{2},\ldots,d_{I};x\right)\right]$ (12)
$+\tau^{-1}\mathbb{E}_{d_{i}\sim\rho\left(d;\pi^{*}\left(x\right)\right)}\Big{[}l\left(d_{1},d_{2},\ldots,d_{I};x\right)$
$\Big{(}\sum_{i}{\sum_{h}{\nabla\left(Q_{s_{h}^{i}a_{h}^{i}}^{*}\left(x\right)-V_{s_{h}^{i}}^{*}\left(x\right)\right)}}\Big{)}\Big{]}.$ (13)
Essentially, the first term (12) in the hyper-gradient contributes to decreasing the upper-level function, and the second term (13) aggregates the gradient information transmitted from the lower level. Assembling these two directions constructs the hyper-gradient, and updating the upper-level variable $x$ along $-\nabla\phi(x)$ will lead to a first-order stationary point (see Theorem 7.6).
@@
~$\widetilde{\nabla}V_{s}(x_{k},\pi_{k})$ $=\mathbb{E}\big{[}\sum_{t=0}^{\infty}\gamma^{t}\nabla r_{s_{t},a_{t}}(x_{k})\,\big{|}\,s_{0}=s,\pi_{k}\big{]},$
~$\widetilde{\nabla}Q_{sa}(x_{k},\pi_{k})$ $=\mathbb{E}\big{[}\sum_{t=0}^{\infty}\gamma^{t}\nabla r_{s_{t},a_{t}}(x_{k})\,\big{|}\,s_{0}\!=\!s,a_{0}\!=\!a,\pi_{k}\big{]};$
@@
~$\widetilde{\nabla}\phi(x_{k},\pi_{k})=$ $\ \mathbb{E}_{d_{i}\sim\rho\left(d;\pi_{k}\right)}\left[\nabla l\left(d_{1},d_{2},\ldots,d_{I};x_{k}\right)\right]$ (14)
$+\tau^{-1}\mathbb{E}_{d_{i}\sim\rho\left(d;\pi_{k}\right)}\Big{[}l\left(d_{1},d_{2},\ldots,d_{I};x_{k}\right)$
$\Big{(}\sum_{i}{\sum_{h}{\widetilde{\nabla}\left(Q_{s_{h}^{i}a_{h}^{i}}-V_{s_{h}^{i}}\right)\left({x_{k},\pi_{k}}\right)}}\Big{)}\Big{]}.$
Consequently, we propose a model-free soft bilevel reinforcement learning algorithm, called SoBiRL, which is outlined in Algorithm 1. In the $k$-th outer iteration, we search for an approximate optimal soft policy $\pi_{k}$ satisfying ${\left\|{{\pi}_{k}-\pi^{*}(x_{k})}\right\|^{2}_{2}\leq\epsilon}$, which can be achieved by executing $\mathcal{O}\left({\log\epsilon^{-1}}\right)$ iterations of the policy mirror descent algorithm (Lan,, 2023; Zhan et al.,, 2023). Then, we utilize ~$\widetilde{\nabla}\phi(x_{k},\pi_{k})$, whose estimation only involves first-order oracles, for an inexact hyper-gradient descent step.
@@
2:     Solve the lower-level problem with the initialization $\pi_{k-1}$ to get ${\pi}_{k}$ with $\left\|{{\pi}_{k}-\pi^{*}_{k}}\right\|_{2}^{2}\leq\epsilon$
3:     Compute ~$\widetilde{\nabla}\phi(x_{k},\pi_{k})$